# Multi-Agent Bandit Learning through Heterogeneous Action Erasure Channels

### v1

@@ I INTRODUCTION
In this work, we dispense with the need for feedback. We ask, what performance can we achieve if the learner action requests are delivered according to a known probabilistic model, but we have no additional information on whether each specific request is delivered or not. In particular, we assume operation over $T$ rounds, where at each round a central learner sends commands (which action to play) to $M$ agents through erasure channels with erasure probability $\epsilon_{i}$, $i=1\ldots M$, where these probabilities can be arbitrarily different across channels. This induces a probability distribution on the reception time of each action request, different for each agent [^1] We note that our schemes extend to more general such probability distributions.. The agents send no feedback (thus the learner does not know which action request the agents are following); the agents play at each round the last action command they received. The learner observes the reward for the (uknown) action played through an error-free channel - which can lead to erroneous action-reward associations. Indeed, energy and space limitations of micro-robots, security constraints in military applications, or simply the structure of the communication protocol can prevent transmitting feedback back to the learner; and agents are required to perform an action at each time, since even no action (for instance staying still) is also an action, see Table I for a small example.
@@
Various MAB algorithms achieving optimal or near-optimal regret bounds under different assumptions have been proposed over the years [10]. Previous studies designed optimal algorithms for the simple MAB setting through providing gap dependent regret of $\tilde{O}(\sum_{a:\Delta_{a}>0}\frac{1}{\Delta_{a}})$ and worst-case regret of $O(\sqrt{KT\log{T}})$ [11, 12, 13] . However, these algorithms are not resilient to action erasures and they are not optimized for multi-agent settings.
@@
Multi-armed bandits with delayed feedback has been studied in recent years under different settings due to its practical applications. For the stochastic setting, [14] shows that bounded unknown i.i.d. delays cause an additive increase in the regret, i.e., $O(\sqrt{KT\log{T}}+K\mathbb{E}\left[D\right])$ where the first term is the regret of stochastic MAB problem with no delays at round $T$, $K$ is the number of actions, and $\mathbb{E}\left[D\right]$ is the expected delay. Following the work in [14], [15] proposes a queue-based MAB algorithm to handle delays. Later, [16] achieves the same additive increase in regret as in [14] under delayed aggregated anonymous feedback. [17] studies Bernoulli bandits with known delay distribution where some feedback could also be censored, i.e., do not reach the learner. Relevant to these works, [18] proposes an algorithm for the best arm identification problem in stochastic MABs with partial and delayed feedback where the aim is to minimize the number of samples for identifying the best action. They extend their methods to the parallel MAB setting, i.e., multiple actions are pulled at each time; however, they provide no lower bounds on the sample complexity of their problem setting. While these works incorporate delays into the stochastic MAB model, delays are associated with action pulls whereas, in our setup,delays are associated with agents instead.
@@
In Section II we introduce the notation and system model; we explain the proposed algorithm in Section III; analyze it in Section IV and provide upper bounds; evaluate and compare with possible baselines in Section V; and conclude in Section VI.
@@ II-B Multi-Agent Multi-Armed Bandits with Action Erasures
Consider a central learner connected to $M$ distributed agents, indexed by $[M]$, over heterogeneous erasure channels. The learner faces a stochastic $K$-armed bandit problem, i.e., $|{\mathcal{A}}|=K$. At each round $t$ during a time horizon $T$, the learner selects an action $a_{t}^{(m)}\in{\mathcal{A}}$ for each agent $m\in[M]$ to play. That is, $M$ actions are played per round (unlike the traditional setting described above). When an action is chosen for agent $m$, it is communicated through an i.i.d. action erasure channel characterized by erasure probability $\epsilon_{m}$, and may or may not be received. That is, independently from other rounds and agents, each agent $m$ receives $a_{t}^{(m)}$ with probability $1-\epsilon_{m}$ and does not receive an action with probability $\epsilon_{m}$. The learner does not know which action requests get erased, but has knowledge of upper bounds on the erasure probabilities. The agents, on the other hand, perceive their own erasures but they do not have a feedback mechanism to inform the learner, i.e., there is no uplink between the agents and the learner. Furthermore, as motivated from applications discussed in Section I, we assume that the agents cannot (or do not wish to) run the algorithm themselves and and continues to play the same action (last successfully received action), denoted as $\tilde{a}_{t}^{(m)}\in{\mathcal{A}}$ to play in the case of an erasure. $\tilde{a}_{0}^{(m)}\in{\mathcal{A}}$ denotes the action performed by the agent $m$ if the action in first round is erased, it is chosen uniformly at random. An example of multi-agent multi-armed bandit learning with action erasures is provided in Table I.
@@ III PROPOSED ALGORITHM
Example 2 Another straightforward approach is to first complete $4^{i}$ pulls for one action by distributing the pulls across all agents, and then move onto the next. That is, for each action, the learner sends it to all agents, and waits until $4^{i}$ effective pulls (i.e., not counting repetitions) are received back. Note that even if an agent $m$ needs to play one pull, we still need to wait first for $\alpha_{m}$ rounds before collecting this effective reward. This scheduling has an end time $T^{(i)}=K^{(i)}\min\limits_{\tilde{M}\in[M]}(\frac{\sum_{m=1}^{\tilde{M}}\alpha_{m}+4^{i}}{\tilde{M}})$, where $K^{(i)}=|{\mathcal{A}}_{i}|$. This can be suboptimal; for instance, if $\alpha_{m}=\alpha~{}\forall m\in[M]$, then the end time of this scheduling is $T^{(i)}=K^{(i)}\alpha+K^{(i)}\lceil\frac{4^{i}}{M}\rceil$, whereas an end time of $\lceil\frac{K^{(i)}}{M}\rceil(\alpha+4^{i})$ can be achieved using the scheduling explained in Example 1.
$\tau:=4^{i}K^{(i)}/\sum_{m=1}^{M}1/(\frac{\alpha_{m}}{4^{i}}+1).$ (1)
@@
Connecting to channels with delays. We note that our algorithm BatchSP2 (and its analysis, in the next section), directly applies to channels with delays, where an action sent by the learner to an agent $m$ is received after $t$ rounds with some probability $p_{t}^{(m)}$, according to some probability distribution. Indeed, although we used erasure channels for our narrative in this paper, and motivated use of repetitions over such channels, the only fact that BatchSP2 essentially hinges on is that, agent $m$ will receive a sent action with probability at least $\frac{1}{T}$ after $\alpha_{m}$ rounds, where $\alpha_{m}$ is known. In our case $\alpha_{m}$ was dictated from the repetition protocol, in other setups it could be dictated from delivery delay or delivery uncertainty.
@@ IV REGRET ANALYSIS
Proof Sketch of Theorem 1. The regret bound is achieved by decomposing the regret of each batch as $\mathbb{E}[R_{T}^{(i)}]=\sum_{a}T_{ia}\Delta_{a}$ and bounding the expected number of times $\mathbb{E}[T_{ia}]$ that arm $a$ is pulled in batch $i$. To that end, we condition on a good event entailing that for each agent $m$ and each consecutive $\alpha_{m}+x$ pulls from action $a$, the last $x$ rewards are samples from the distribution of action $a$. This provides a concentration of the empirical means used in Algorithm 1 with high probability. As a result, we get an upper bound on the number of batches a suboptimal arm can survive. Having this, to bound $T_{ia}$, it only remains to bound the number of times an active action is pulled in batch $i$. This is proved by utilizing the upper bound on the number of pulls in Lemma 1 and using the symmetry imposed by the randomization in Algorithm 2 to show that each action has an equal contribution in the total number of pulls. The final regret bound is obtained by showing that the good event has high probability.
@@
where Δsubscriptnormal-$\Delta_{\star}$ is the value satisfying
@@
Proof Sketch of Theorem 2. The regret bound is proved by bounding the regret bound for arms with small gap (less than $\Delta_{\star}$) by $TM\Delta_{\star}$ and using the bound in Theorem 1 for the remaining arms with large gap (greater than $\Delta_{\star}$). The value of $\Delta_{\star}$ is chosen to balance the regret resulting from arms with small gaps and the regret from arms with large gaps.
@@ V EXPERIMENTS
We have explored a number of experimental setups (in terms of number of actions, channel quality, horizons, etc); we here show results for two experiments that we believe are representative:
@@ Appendix A LINEAR PROGRAM FORMULATION
where ${\bm{X}}\in\mathbb{R}^{M\times K}$ captures the variables of the program, with $X_{mk}$ indicating the number of effective pulls of action $k$ performed by agent $m$. The objective function (2a) is to minimize the latest end time among agents. Constraint (2b) ensures that the total number of effective pulls for each action is $4^{i}$; and constraint (2c) forces effective pulls assigned to each agent per action to be an integral value in $[0,4^{i}]$. It is easy to see that the progam in (A) is equivalent to the following integer linear program (ILP):
$\min\limits_{\begin{subarray}{c}{\bm{X}},{\bm{W}}\in\mathbb{R}^{M\times K}\\ t\in\mathbb{R}\end{subarray}}$ $~{}~{}~{}~{}t$ (3c)
s.t. $~{}~{}\sum\limits_{k=1}^{K}\left(\alpha_{m}W_{mk}+X_{mk}\right)\leq t,~{}~{}~{}\forall m\in[M]$ (3d)
$~{}~{}X_{mk}\leq 4^{i}W_{mk}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (3e)
$~{}~{}\sum\limits_{m=1}^{M}X_{mk}=4^{i}~{}~{}~{}\forall k\in[K]$ (3f)
$~{}~{}X_{mk}\in\{0,1,2,\cdots,4^{i}\}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (3g)
$~{}~{}W_{mk}\in\{0,1\}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K],$ (3h)
where the variable $t\in\mathbb{R}$ replaces the $\max$ in objective equation 2a and the variable ${\bm{W}}\in\mathbb{R}^{M\times K}$ replaces the indicator function. Notice that for any feasible solution ${\bm{X}}$, if $X_{mk}>0$, $W_{mk}=1$. The relaxed version of the ILP in (A) can be written as
$\min\limits_{\begin{subarray}{c}{\bm{X}},{\bm{W}}\in\mathbb{R}^{M\times K}\\ t\in\mathbb{R}\end{subarray}}$ $~{}~{}~{}~{}t$ (4c)
s.t. $~{}~{}\sum\limits_{k=1}^{K}\left(\alpha_{m}W_{mk}+X_{mk}\right)\leq t,~{}~{}~{}\forall m\in[M]$ (4d)
$~{}~{}{\bm{X}}_{mk}\leq 4^{i}W_{mk}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (4e)
$~{}~{}\sum\limits_{m=1}^{M}X_{mk}=4^{i}~{}~{}~{}\forall k\in[K]$ (4f)
$~{}~{}0\leq X_{mk}\leq 4^{i}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (4g)
$~{}~{}0\leq W_{mk}\leq 1~{}~{}~{}\forall m\in[M],~{}\forall k\in[K],$ (4h)
Notice that the minimum value $W_{mk}$ can take is $X_{mk}/4^{i}$ due to (4e); hence, by replacing $W_{mk}$ with its minimum value, we get the following linear program which gives a lower bound on the ILP (A):
$\min\limits_{\begin{subarray}{c}{\bm{X}}\in\mathbb{R}^{M\times K}\\ t\in\mathbb{R}\end{subarray}}$ $~{}~{}~{}~{}t$ (5c)
s.t. $~{}~{}\sum\limits_{k=1}^{K}X_{mk}\left(\frac{\alpha_{m}}{4^{i}}+1\right)\leq t~{}~{}~{}\forall m\in[M]$ (5d)
$~{}~{}\sum\limits_{m=1}^{M}X_{mk}=4^{i}~{}~{}~{}\forall k\in[K]$ (5e)
$~{}~{}0\leq X_{mk}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K],$ (5f)
In the linear program (A), $X_{mk}$ is the variable that indicates how many effective pulls are assigned to agent $m$ for action $k$. (5e) forces each action to be pulled $4^{i}$ effective times; however, instead of an integer number of pulls, each agent is allowed to perform nonnegative fractional pulls. Furthermore, (5d) indicates that for each agent $m$, one effective pull has a cost of $\frac{\alpha_{m}}{4^{i}}+1$.
@@
The optimal objective value of (A) satisfies, $t^{\star}=\sum_{k=1}^{K}X_{mk}^{\star}(\alpha_{m}/4^{i}+1)~{}\forall m\in M$, where ($t^{\star},{\bm{X}}^{\star}$) is the optimal solution of A.
Proof of Claim 1. First, we observe that at least one of the inequalities in (5d) holds with equality, otherwise the value of $t^{\star}$ can be decreased leading to a better objective. Define the set of indices ${\mathcal{E}}_{i}:=\{m\in[M]:\sum_{k=1}^{K}X_{mk}^{\star}(\alpha_{m}/4^{i}+1)=t^{\star}\}$.
@@
forms a feasible solution in (A) with a smaller objective value $t^{\prime}<t^{\star}$; hence, $(t^{\star},{\bm{X}}^{\star})$ cannot be optimal. Then at the optimal solution $(t^{\star},{\bm{X}}^{\star})$, $t^{\star}=\sum_{k=1}^{K}X_{mk}^{\star}(\alpha_{m}/4^{i}+1)~{}\forall m\in M$.
Using Claim 1 and the constraint (5e);
@@
Note that the solution of the relaxed LP (A) can be directly used for scheduling of actions by adding $\max{(2\alpha_{M-1},\alpha_{m})}$ to the end time $t^{\star}$. Since the relaxation in general does not give a feasible solution for the ILP, we add $\max{(2\alpha_{M-1},\alpha_{M})}$ to the end time of the relaxed ILP to guarantee a feasible solution for the ILP. As the additional time slots accumulate regret across all agents, this can result in $\Omega(M\alpha_{M})$ additional regret which can be large for large $M$. Our algorithm improves the $M$ factor in $M\alpha_{M}$.

### v2

@@ I INTRODUCTION
In this work, we dispense with the need for feedback. We ask, what performance can we achieve if the learner action requests are delivered according to a known probabilistic model, but we have no additional information on whether each specific request is delivered or not. In particular, we assume operation over $T$ rounds, where at each round a central learner sends commands (which action to play) to $M$ agents through erasure channels with erasure probability $\epsilon_{i}$, $i=1\ldots M$, where these probabilities can be arbitrarily different across channels. This induces a Geometric distribution on the reception time of each action request, different for each agent [^1] We note that our schemes extend to more general such probability distributions; see Section III-A for a detailed discussion.. The agents send no feedback (thus the learner does not know which action request the agents are following); the agents play at each round the last action command they received. The learner observes the reward for the (uknown) action played through an error-free channel - which can lead to erroneous action-reward associations. Indeed, energy and space limitations of micro-robots, security constraints in military applications, or simply the structure of the communication protocol can prevent transmitting feedback back to the learner; and agents are required to perform an action at each time, since even no action (for instance staying still) is also an action, see Table I for a small example.
@@
Various MAB algorithms achieving optimal or near-optimal regret bounds under different assumptions have been proposed over the years [10]. Previous studies designed optimal algorithms for the simple MAB setting through providing gap dependent regret of $\tilde{O}(\sum_{a:\Delta_{a}>0}\frac{1}{\Delta_{a}})$ and worst-case regret of $O(\sqrt{KT\log{T}})$ [11, 12, 13] . However, these algorithms are not resilient to action erasures and they are not optimized for multi-agent settings. Building upon MAB algorithms, there has been a considerable amount of recent work on multi-agent MABs in various settings [14, 15, 16, 17]. However, these works predominantly consider connected agents with some form of communication between neighbors or improve the regret bounds through feedback mechanisms, primarily involving collision sensing for agents [18, 19]. Consequently, they fail to effectively handle action erasures, particularly when connections between agents and feedback mechanism are missing.
@@
Multi-armed bandits with delayed feedback has been studied in recent years under different settings due to its practical applications. For the stochastic setting, [20] shows that bounded unknown i.i.d. delays cause an additive increase in the regret, i.e., $O(\sqrt{KT\log{T}}+K\mathbb{E}\left[D\right])$ where the first term is the regret of stochastic MAB problem with no delays at round $T$, $K$ is the number of actions, and $\mathbb{E}\left[D\right]$ is the expected delay. Following the work in [20], [21] proposes a queue-based MAB algorithm to handle delays. Later, [22] achieves the same additive increase in regret as in [20] under delayed aggregated anonymous feedback. [23] studies Bernoulli bandits with known delay distribution where some feedback could also be censored, i.e., do not reach the learner. Relevant to these works, [24] proposes an algorithm for the best arm identification problem in stochastic MABs with partial and delayed feedback where the aim is to minimize the number of samples for identifying the best action. They extend their methods to the parallel MAB setting, i.e., multiple actions are pulled at each time; however, they provide no lower bounds on the sample complexity of their problem setting. While these works incorporate delays into the stochastic MAB model, delays are associated with action pulls whereas, in our setup, delays are associated with agents and are independent on the pulled action for the same agent.
@@
In Section II, we introduce the notation and system model; we explain the proposed algorithm in Section III; analyze it in Section IV and provide upper bounds; evaluate and compare with possible baselines in Section V; and conclude in Section VI.
@@ II-B Multi-Agent Multi-Armed Bandits with Action Erasures
Consider a central learner connected to $M$ distributed agents, indexed by $[M]$, over heterogeneous erasure channels. The learner faces a stochastic $K$-armed bandit problem, i.e., $|{\mathcal{A}}|=K$. At each round $t$ during a time horizon $T$, the learner selects an action $a_{t}^{(m)}\in{\mathcal{A}}$ for each agent $m\in[M]$ to play. That is, $M$ actions are played per round (unlike the traditional setting described above). When an action is chosen for agent $m$, it is communicated through an i.i.d. action erasure channel characterized by erasure probability $\epsilon_{m}$, and may or may not be received. That is, independently from other rounds and agents, each agent $m$ receives $a_{t}^{(m)}$ with probability $1-\epsilon_{m}$ and does not receive an action with probability $\epsilon_{m}$. The learner does not know which action requests get erased, but has knowledge of upper bounds on the erasure probabilities. The agents, on the other hand, perceive their own erasures but they do not have a feedback mechanism to inform the learner, i.e., there is no uplink between the agents and the learner. Furthermore, as motivated from applications discussed in Section I, we assume that the agents cannot (or do not wish to) run the algorithm themselves and continues to play the same action (last successfully received action), denoted as $\tilde{a}_{t}^{(m)}\in{\mathcal{A}}$ to play in the case of an erasure. $\tilde{a}_{0}^{(m)}\in{\mathcal{A}}$ denotes the action performed by the agent $m$ if the action in first round is erased, it is chosen uniformly at random. An example of multi-agent multi-armed bandit learning with action erasures is provided in Table I.
@@ III PROPOSED ALGORITHM
Example 2 Another straightforward approach is to first complete $4^{i}$ pulls for one action by distributing the pulls across all agents, and then move onto the next. That is, for each action, the learner sends it to all agents, and waits until $4^{i}$ effective pulls (i.e., not counting repetitions) are received back. Note that even if an agent $m$ needs to play one pull, we still need to wait first for $\alpha_{m}$ rounds before collecting this effective reward. This scheduling has an end time $T^{(i)}=K^{(i)}\min\limits_{\tilde{M}\in[M]}(\frac{\sum_{m=1}^{\tilde{M}}\alpha_{m}+4^{i}}{\tilde{M}})$, where $K^{(i)}=|{\mathcal{A}}_{i}|$. This can be suboptimal, e.g., if $\alpha_{m}=\alpha~{}\forall m\in[M]$, then the end time of this scheduling is $T^{(i)}=K^{(i)}\alpha+K^{(i)}\lceil\frac{4^{i}}{M}\rceil$, whereas an end time of $\lceil\frac{K^{(i)}}{M}\rceil(\alpha+4^{i})$ can be achieved using the scheduling explained in Example 1.
@@
III-A Connecting to Channels with Delays
We note that our algorithm BatchSP2 (and its analysis, in the next section), directly applies to channels with delays, where an action sent by the learner to an agent $m$ is received after $t$ rounds with some probability $p_{t}^{(m)}$. Indeed, although we used erasure channels for our narrative in this paper, and motivated use of repetitions over such channels, the only fact that BatchSP2 essentially hinges on is that, agent $m$ will receive a sent action with probability at least $\frac{1}{T}$ after $\alpha_{m}$ rounds, where $\alpha_{m}$ is known. It implies that any known (or estimated) delay/probability of successful reception can be used with our algorithms. In our case $\alpha_{m}$ was dictated from the repetition protocol, in other setups it could be dictated from delivery delay or delivery uncertainty. Datasets and models in literature, e.g., [27] or [28], can provide empirical values for delay/probability in such setups.
@@ IV REGRET ANALYSIS
Proof Sketch of Theorem 1. The regret bound is achieved by decomposing the regret of each batch as $\mathbb{E}[R_{T}^{(i)}]=\sum_{a}T_{ia}\Delta_{a}$ and bounding the expected number of times $\mathbb{E}[T_{ia}]$ that arm $a$ is pulled in batch $i$. To that end, we condition on a good event entailing that for each agent $m$ and each consecutive $\alpha_{m}+x$ pulls from action $a$, the last $x$ rewards are samples from the distribution of action $a$. This provides a concentration of the empirical means used in Algorithm 1 with high probability. As a result, we get an upper bound on the number of batches a suboptimal arm can survive. Having this, to bound $T_{ia}$, it only remains to bound the number of times an active action is pulled in batch $i$, which is highly sensitive to the scheduling of action pulls. This is proved by utilizing the upper bound on the number of pulls in Lemma 1 and leveraging the symmetry imposed by the randomization in Algorithm 2 to show that each action has an equal contribution in the total number of pulls. The final regret bound is obtained by showing that the good event has high probability.
@@
The three components of the regret bound in Theorem 1 originate from distinct aspects of the algorithm. The initial term, $\sum_{a:\Delta_{a}>0}\log{(KMT)}/\Delta_{a}$, is inevitable, representing the order optimal regret achievable under perfect channels (no delay, no erasure). The second and third terms are due to the repetition and scheduling of actions (Algorithm 2). It is noteworthy that, the second term matches the regret of an optimal scheduling algorithm (see App. A). Additionally, under perfect channels, this term simplifies to the a lower bound on the regret up to logarithmic factors. The third term, $\sum_{m=1}^{M}\alpha_{m}\log{(MT)}$, emerges at the outset of the learning process, reflecting that each agent $m$ will repeat the first pulled (suboptimal) action for $\alpha_{m}$ iterations on average.
@@
where $\Delta_{\star}$ is the value satisfying
@@
Proof Sketch of Theorem 2. The regret bound is proved by bounding the regret bound for arms with small gap (less than $\Delta_{\star}$, which will be determined later) by $TM\Delta_{\star}$ and using the bound in Theorem 1 for the remaining arms with large gap (greater than $\Delta_{\star}$). The value of $\Delta_{\star}$ is chosen to minimize the bound by balancing the regret resulting from arms with small gaps and the regret from arms with large gaps.
@@ V EXPERIMENTS
We have explored a number of experimental setups (in terms of number of actions, channel quality, horizons, etc) [^5] The code to our experiments is available here.; we here show results for two experiments that we believe are representative:
@@ VI CONCLUSION
Acknowledgements
We thank the anonymous reviewers and the meta-reviewer for their insightful suggestions and comments. This work is supported in part by NSF grants #2007714 and #2221871, by Army Research Laboratory grant under Cooperative Agreement W911NF-17-2-0196, and by the Amazon Faculty Award.
@@ Appendix A LINEAR PROGRAM FORMULATION
where ${\bm{X}}\in\mathbb{R}^{M\times K}$ captures the variables of the program, with $X_{mk}$ indicating the number of effective pulls of action $k$ performed by agent $m$. The objective function (2a) is to minimize the latest end time among agents. Constraint (2b) ensures that the total number of effective pulls for each action is $4^{i}$; and constraint (2c) forces effective pulls assigned to each agent per action to be an integral value in $[0,4^{i}]$. It is easy to see that the progam in (2) is equivalent to the following integer linear program (ILP):
$\min\limits_{\begin{subarray}{c}{\bm{X}},{\bm{W}}\in\mathbb{R}^{M\times K}\\ t\in\mathbb{R}\end{subarray}}$ $~{}~{}~{}~{}t$ (3a)
s.t. $~{}~{}\sum\limits_{k=1}^{K}\left(\alpha_{m}W_{mk}+X_{mk}\right)\leq t,~{}~{}~{}\forall m\in[M]$ (3b)
$~{}~{}X_{mk}\leq 4^{i}W_{mk}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (3c)
$~{}~{}\sum\limits_{m=1}^{M}X_{mk}=4^{i}~{}~{}~{}\forall k\in[K]$ (3d)
$~{}~{}X_{mk}\in\{0,1,2,\cdots,4^{i}\}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (3e)
$~{}~{}W_{mk}\in\{0,1\}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K],$ (3f)
where the variable $t\in\mathbb{R}$ replaces the $\max$ in objective equation 2a and the variable ${\bm{W}}\in\mathbb{R}^{M\times K}$ replaces the indicator function. Notice that for any feasible solution ${\bm{X}}$, if $X_{mk}>0$, $W_{mk}=1$. The relaxed version of the ILP in (3) can be written as
$\min\limits_{\begin{subarray}{c}{\bm{X}},{\bm{W}}\in\mathbb{R}^{M\times K}\\ t\in\mathbb{R}\end{subarray}}$ $~{}~{}~{}~{}t$ (4a)
s.t. $~{}~{}\sum\limits_{k=1}^{K}\left(\alpha_{m}W_{mk}+X_{mk}\right)\leq t,~{}~{}~{}\forall m\in[M]$ (4b)
$~{}~{}{\bm{X}}_{mk}\leq 4^{i}W_{mk}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (4c)
$~{}~{}\sum\limits_{m=1}^{M}X_{mk}=4^{i}~{}~{}~{}\forall k\in[K]$ (4d)
$~{}~{}0\leq X_{mk}\leq 4^{i}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K]$ (4e)
$~{}~{}0\leq W_{mk}\leq 1~{}~{}~{}\forall m\in[M],~{}\forall k\in[K],$ (4f)
Notice that the minimum value $W_{mk}$ can take is $X_{mk}/4^{i}$ due to (4c); hence, by replacing $W_{mk}$ with its minimum value, we get the following linear program which gives a lower bound on the ILP (3):
$\min\limits_{\begin{subarray}{c}{\bm{X}}\in\mathbb{R}^{M\times K}\\ t\in\mathbb{R}\end{subarray}}$ $~{}~{}~{}~{}t$ (5a)
s.t. $~{}~{}\sum\limits_{k=1}^{K}X_{mk}\left(\frac{\alpha_{m}}{4^{i}}+1\right)\leq t~{}~{}~{}\forall m\in[M]$ (5b)
$~{}~{}\sum\limits_{m=1}^{M}X_{mk}=4^{i}~{}~{}~{}\forall k\in[K]$ (5c)
$~{}~{}0\leq X_{mk}~{}~{}~{}\forall m\in[M],~{}\forall k\in[K],$ (5d)
In the linear program (5), $X_{mk}$ is the variable that indicates how many effective pulls are assigned to agent $m$ for action $k$. (5c) forces each action to be pulled $4^{i}$ effective times; however, instead of an integer number of pulls, each agent is allowed to perform nonnegative fractional pulls. Furthermore, (5b) indicates that for each agent $m$, one effective pull has a cost of $\frac{\alpha_{m}}{4^{i}}+1$.
@@
The optimal objective value of (5) satisfies, $t^{\star}=\sum_{k=1}^{K}X_{mk}^{\star}(\alpha_{m}/4^{i}+1)~{}\forall m\in M$, where ($t^{\star},{\bm{X}}^{\star}$) is the optimal solution of 5.
Proof of Claim 1. First, we observe that at least one of the inequalities in (5b) holds with equality, otherwise the value of $t^{\star}$ can be decreased leading to a better objective. Define the set of indices ${\mathcal{E}}_{i}:=\{m\in[M]:\sum_{k=1}^{K}X_{mk}^{\star}(\alpha_{m}/4^{i}+1)=t^{\star}\}$.
@@
forms a feasible solution in (5) with a smaller objective value $t^{\prime}<t^{\star}$; hence, $(t^{\star},{\bm{X}}^{\star})$ cannot be optimal. Then at the optimal solution $(t^{\star},{\bm{X}}^{\star})$, $t^{\star}=\sum_{k=1}^{K}X_{mk}^{\star}(\alpha_{m}/4^{i}+1)~{}\forall m\in M$.
Using Claim 1 and the constraint (5c);
@@
Note that the solution of the relaxed LP (5) can be directly used for scheduling of actions by adding $\max{(2\alpha_{M-1},\alpha_{m})}$ to the end time $t^{\star}$. Since the relaxation in general does not give a feasible solution for the ILP, we add $\max{(2\alpha_{M-1},\alpha_{M})}$ to the end time of the relaxed ILP to guarantee a feasible solution for the ILP. As the additional time slots accumulate regret across all agents, this can result in $\Omega(M\alpha_{M})$ additional regret which can be large for large $M$. Our algorithm improves the $M$ factor in $M\alpha_{M}$.